a man of mass 50 kg is pulling on a plank of mass 100 kg kept on a smooth floor as shown with force of 100 N.  if both man & plank move togeather, find force of friction acting on man

Grade:11

2 Answers

Smit Dama
6 years ago
Draw free body diagram for the man and plank.The man is standing on the plank. Then he is pulling the rope towards himself. He pushes the plank forward with the legs so that he can pull the rope backwards. So the friction f acts in the forward direction for the plank. Also f acts in the opposite direction on the man. The tension T in the rope is 100 N (given). Let the friction force = f Newtons.Let the common acceleration = a m/s^2Man: net force = T - f = m a = 50 a Plank: net force = T + f = m a = 100 aAs T = 100 N, a = 4/3 m/s^2 and f = 100/3 Newtons.
Raj pauwaa
5 years ago
Draw free body diagram for the man and plank.
The man is standing on the plank. Then he is pulling the rope towards himself. He pushes the plank forward with the legs so that he can pull the rope backwards.
   So the friction f acts in the forward direction for the plank. Also f acts in the opposite direction on the man.  The tension T in the rope is 100 N (given). Let the friction force = f Newtons.
Let the common acceleration = a  m/s^2
Man:    net force = T - f = m a = 50 a 
Plank:  net force = T + f = m a = 100 a
As T = 100 N,     a = 4/3 m/s^2     and   f = 100/3 Newtons.